OLIVER69:

Help Please Divide using long division. Show work (9x^2 -21x - 20) ÷ (x - 1)

Arieonna:

@oliver69

Arieonna:

\[\frac{ (9x²-21x-20) }{( x-1) }\]

Arieonna:

\[= 9x+\frac{ -12x-20 }{ x-1}\]

Arieonna:

\[= 9x = 12+\frac{ -32 }{ x - 1 }\]

Arieonna:

\[= 9x-12-\frac{ 32 }{ x-1 }\]

Arieonna:

hope that helped u I tried my best explaining it

OLIVER69:

@arieonna wrote:
I hope that helped you I tried my best to explain it
thanks by the way, sorry I wasn't on to help you with your stuff

Arieonna:

its ok i did most of my work yesturday with out help and got most of them correct doing it by my self